OpenStudy (anonymous):

how can the radian measure of an angle determine the arc length on the unit circle?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

what's the radius on a unit circle?

OpenStudy (anonymous):

idu your question

OpenStudy (jdoe0001):

well.. do you know what a unit circle is?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

so.... any ideas on how long its radius is?

OpenStudy (anonymous):

2pi

OpenStudy (anonymous):

wait no 1 sorry

OpenStudy (jdoe0001):

well... recall the UNIT circle, has a unit radius, thus the "unit circle" 

OpenStudy (anonymous):

yep

OpenStudy (jdoe0001):

\(\bf \textit{arc's length}=s={\color{red}{ r}}\theta\impliedby \textit{in radians}\qquad{\color{red}{ r}}=1\qquad s={\color{red}{ 1}}\theta \\ \quad \\ s=\theta\)

OpenStudy (anonymous):

now you lost me